OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a18 = 97, a20 = 281 an = -1467 - 92(n + 1) an = -1467 + 92(n + 1) an = -1467 + 92(n - 1) an = -1467 + 92(n - 2)

OpenStudy (anonymous):

the expression for nth term of an A.p. is : tn= a + (n-1)d thus, a+17d=97 and a+19d=281 solvin these eqns....a=-1467 and d= 92 put these values in the eqn...

OpenStudy (anonymous):

so this one an = -1467 + 92(n - 1) ??

OpenStudy (anonymous):

yup